### v1

# The Boolean spectrum of a Grothendieck category

Abstract.
A notion of support for objects in any Grothendieck category is introduced. This is based on the spectral category of a Grothendieck category and uses its Boolean lattice of localising subcategories. The support provides a classification of all subcategories that are closed under arbitrary coproducts, subobjects, and essential extensions. There is also a notion of exact support which classifies certain thick subcategories. As an application we extend for any ring Crawley-Boevey’s correspondence between definable subcategories of modules and closed subsets of the Ziegler spectrum.
Key words and phrases:
Definable subcategory, Grothendieck category, spectral category, support, Ziegler spectrum
2020 Mathematics Subject Classification:
18E10 (primary), 16D70, 18E40, 18E45 (secondary)
1. Introduction
The spectrum $\operatorname{Sp}\mathcal{A}$ of a Grothendieck category $\mathcal{A}$ is the set of isomorphism classes of indecomposable injective objects. This definition is due to Gabriel, and he uses this spectrum when he reconstructs a noetherian scheme from its category of quasi-coherent sheaves [10]. A finer invariant is introduced in work of Gabriel and Oberst [11]. The spectral category $\operatorname{Spec}\mathcal{A}$ is again a Grothendieck category with the additional propery that every short exact sequence is split exact. There is a canonical functor $\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ which identifies $\operatorname{Sp}\mathcal{A}$ with the isomorphism classes of simple objects in $\operatorname{Spec}\mathcal{A}$.
In this work we study the lattice $\mathbf{L}(\operatorname{Spec}\mathcal{A})$ of localising subcategories of $\operatorname{Spec}\mathcal{A}$, which one may think of as an intermediate between $\operatorname{Sp}\mathcal{A}$ and $\operatorname{Spec}\mathcal{A}$. This is actually a complete Boolean lattice; so it identifies via Stone duality with the Boolean lattice of clopen sets of its spectrum
$\operatorname{Spc}\mathcal{A}:=\operatorname{Spec}(\mathbf{L}(\operatorname{Spec}\mathcal{A}))$
which we call the Boolean spectrum of $\mathcal{A}$. There is a canonical embedding
$\operatorname{Sp}\mathcal{A}\lhook\joinrel\xrightarrow{\ \ \ }\operatorname{Spc}\mathcal{A}$
which is a bijection if and only if $\operatorname{Spec}\mathcal{A}$ is discrete, which means that every object decomposes into a coproduct of simple objects. This happens, for example, when $\mathcal{A}$ is locally noetherian.
To each object $X\in\mathcal{A}$ we assign its support $\operatorname{supp}(X)\subseteq\operatorname{Spc}\mathcal{A}$, and this sets up a bijection
(1.1) $\mathcal{A}\supseteq\mathcal{C}\longmapsto\operatorname{supp}(\mathcal{C})=\bigcup_{X\in\mathcal{C}}\operatorname{supp}(X)$
between the essentially closed subcategories of $\mathcal{A}$ and the clopen subsets of $\operatorname{Spc}\mathcal{A}$ (Theorem 3.3), where a full subcategory is called essentially closed if it is closed under arbitrary coproducts, subobjects, and essential extensions. Let $\langle X\rangle$ denote the smallest essentially closed subcategory containing $X$. The support enjoys the following universal property, which is the analogue of a property of support for tensor triangulated categories due to Balmer [1]. We do not assume any monoidal structure, but our analogue seems reasonable given that one has the identity $\langle X\otimes Y\rangle=\langle X\rangle\cap\langle Y\rangle$ in the tensor triangulated setting.
Theorem (Theorem 3.10).
The map $\operatorname{supp}\colon\operatorname{Ob}\mathcal{A}\to\operatorname{Clop}(\operatorname{Spc}\mathcal{A})$ has the following properties:
  1. (S00)
    $\operatorname{supp}(X)=\varnothing$ for $X=0$.
  2. (S1111)
    $\operatorname{supp}(X)=\operatorname{Spc}\mathcal{A}$ for every generator $X$ of $\mathcal{A}$.
[…]
    $\operatorname{supp}(X)\cup\operatorname{supp}(Y)=\operatorname{supp}(X\oplus Y)$ for all objects $X,Y$.
[…]
    $\operatorname{supp}(X)\cap\operatorname{supp}(Y)=\operatorname{supp}(Z)$ for all objects $X,Y,Z$ with $\langle X\rangle\cap\langle Y\rangle=\langle Z\rangle$.
Moreover, for any pair $(T,\sigma)$ which consists of a topological space $T$ and a map $\sigma\colon\operatorname{Ob}\mathcal{A}\to\operatorname{Clop}(T)$ satisfying the analogue of the above conditions, there exists a unique continuous map $f\colon T\to\operatorname{Spc}\mathcal{A}$ such that $\sigma(X)=f^{-1}(\operatorname{supp}(X))$ for all $X$.
A useful property of this support is that $\operatorname{supp}(X)=\varnothing$ implies $X=0$. This one cannot expect when using $\operatorname{Sp}\mathcal{A}$ as target for a support function, because $\operatorname{Sp}\mathcal{A}$ may be empty. Another option for defining support would be to use as target the frame $\mathbf{L}(\mathcal{A})$ of localising subcategories of $\mathcal{A}$. However, its set of points may be empty as well [34]. The following result shows that $\operatorname{Spc}\mathcal{A}$ is an invariant which encompasses both $\operatorname{Sp}\mathcal{A}$ and $\mathbf{L}(\mathcal{A})$.
Theorem (Theorem 3.18).
Taking a localising subcategory to its support induces an embedding of frames
$\mathbf{L}(\mathcal{A})\hookrightarrow\mathbf{L}(\operatorname{Spec}\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Clop}(\operatorname{Spc}\mathcal{A}).$
For example, when $\mathcal{A}=\operatorname{Mod}\Lambda$ equals the module category of a commutative noetherian ring $\Lambda$, then $\operatorname{Sp}\mathcal{A}\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Spc}\mathcal{A}$ identifies with the set of prime ideals $\operatorname{Spec}\Lambda$, and the above correspondence (1.1) extends Gabriel’s bijection between the localising subcategories of $\operatorname{Mod}\Lambda$ and the specialisation closed subsets of $\operatorname{Spec}\Lambda$, which is one of the ingredients for the reconstruction of a noetherian scheme [10].
The work of Gabriel and his use of the spectrum of indecomposable injectives has been a great inspiration for noncommutative algebraic geometry; see for example [2, 19, 32] where several possible definitions of a spectrum of an abelian category are discussed. Here we follow a somewhat different route. For not necessarily commutative rings there is a spectrum which receives much attention and provides another motivation for this work.
The Ziegler spectrum of a ring $\Lambda$ is given by the isomorphism classes of indecomposable pure-injective $\Lambda$-modules, together with a topology introduced by Ziegler in model-theoretic terms [37]. Let $\operatorname{Ind}\Lambda$ denote the Ziegler spectrum of $\Lambda$. This identifies with the spectrum of a locally coherent Grothendieck category; so the above theory applies. To be more precise, the module category $\operatorname{Mod}\Lambda$ embeds into its purity category $\mathbf{P}(\operatorname{Mod}\Lambda)$, and then $\operatorname{Ind}\Lambda$ identifies with $\operatorname{Sp}\mathbf{P}(\operatorname{Mod}\Lambda)$. The Boolean spectrum provides an extension
(1.2) $\operatorname{Ind}\Lambda\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Sp}\mathbf{P}(\operatorname{Mod}\Lambda)\hookrightarrow\operatorname{Spc}\mathbf{P}(\operatorname{Mod}\Lambda)=:\operatorname{PSpc}\Lambda$
and we propose to study this as an invariant of $\Lambda$.
The Ziegler spectrum has been used in various contexts, for instance when studying representations of Artin algebras, but also in more general categorical settings, including derived and stable categories; see [29] for an encyclopedic survey. An important feature of this spectrum is Crawley-Boevey’s fundamental correspondence
$\mathcal{C}\longmapsto\mathcal{C}\cap\operatorname{Ind}\Lambda$
between definable subcategories of $\operatorname{Mod}\Lambda$ and Ziegler closed subsets of $\operatorname{Ind}\Lambda$; see [5]. In order to illustrate the use of the Boolean spectrum $\operatorname{PSpc}\Lambda$ we extend this correspondence as follows.
Theorem (Theorem 5.6).
The assignment $\mathcal{C}\mapsto\operatorname{supp}(\mathcal{C})$ induces an inclusion preserving bijection between
  1. (1)
    the set of subcategories of $\operatorname{Mod}\Lambda$ that are closed under arbitrary direct sums, pure subobjects, and pure-injective envelopes, and
  2. (2)
    the clopen subsets of $\operatorname{PSpc}\Lambda$.
In particular, the embedding (1.2) identifies $\mathcal{C}\cap\operatorname{Ind}\Lambda$ with $\operatorname{supp}(\mathcal{C})\cap\operatorname{Ind}\Lambda$ when $\mathcal{C}$ is definable.
Returning to the above notion of support $X\mapsto\operatorname{supp}(X)$ for an arbitrary Grothendieck category, we mention that there is also an exact version $X\mapsto\operatorname{supp}_{\mathrm{ex}}(X)$ such that for each exact sequence $0\to X_{1}\to X_{2}\to X_{3}\to 0$ one has
$\operatorname{supp}_{\mathrm{ex}}(X_{i})\subseteq\operatorname{supp}_{\mathrm{ex}}(X_{j})\cup\operatorname{supp}_{\mathrm{ex}}(X_{k})\qquad\text{when}\qquad\{i,j,k\}=\{1,2,3\}.$
This is in line with Neeman’s celebrated classification of localising subcategories for the derived category of a commutative noetherian ring [27], because it agrees with his notion of support for a complex when specialised to a module. The exact support provides a classification of certain thick subcategories (Theorem 4.5), which is an analogue of the bijective correspondence (1.1) and again not requiring any noetherianess assumption. This generalises a classification for modules over a commutative noetherian ring from [35] and another more recent classification for locally noetherian categories [36].
Let us give a quick outline of this paper. In §2 we recall the basic facts about spectral categories and introduce the Boolean spectrum of a Grothendieck category. In §3 we set up the support for Grothendieck categories and establish the classification of essentially closed subcategories. The following §4 provides the derived analogue of the results from §3 which leads to the exact support. The final §5 discusses the applications for module categories, but in the more general setting of exactly definable categories.
2. Spectral Grothendieck categories
Let $\mathcal{A}$ be a Grothendieck category. We recall the notion of its spectral category from the work of Gabriel and Oberst [11]. Then we study the lattice of localising subcategories and use this to define a space, which we call the Boolean spectrum of $\mathcal{A}$.
Spectral categories
We write $\operatorname{Spec}\mathcal{A}$ for the spectral category of $\mathcal{A}$ which is obtained from $\mathcal{A}$ by formally inverting all essential monomorphisms; see [11] or [33, V.7]. The category $\operatorname{Spec}\mathcal{A}$ is again a Grothendieck category in which all exact sequences are split exact. A Grothendieck category with this property is called spectral. The canonical functor
$P\colon\mathcal{A}\longrightarrow\operatorname{Spec}\mathcal{A}:=\mathcal{A}[\mathrm{Ess}^{-1}]$
maps any injective enevelope $X\to E(X)$ in $\mathcal{A}$ to an isomorphism and it identifies the indecomposable injective objects in $\mathcal{A}$ with the simple objects in $\operatorname{Spec}\mathcal{A}$. These properties are clear from the definition. Also, $P$ is left exact and preserves all coproducts since the class of essential monomorphisms is closed under all coproducts. More precisely, for any directed union $X=\sum_{\alpha}X_{\alpha}$ in $\mathcal{A}$ we have $P(X)=\sum_{\alpha}P(X_{\alpha})$. We write $\operatorname{Sp}\mathcal{A}$ for the set of isomorphism classes of indecomposable injective objects in $\mathcal{A}$, and $P$ induces a bijection $\operatorname{Sp}\mathcal{A}\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Sp}(\operatorname{Spec}\mathcal{A})$.
Let $\operatorname{Inj}\mathcal{A}$ denote the full subcategory of injective objects in $\mathcal{A}$. For later use we note that $P$ induces an equivalence
(2.1) $(\operatorname{Inj}\mathcal{A})/\operatorname{Rad}(\operatorname{Inj}\mathcal{A})\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\operatorname{Spec}\mathcal{A},$
where $\operatorname{Rad}\mathcal{C}$ denotes the Jacobson radical of an additive category $\mathcal{C}$; see [24, Proposition 2.5.9]. Recall that $\operatorname{Rad}\mathcal{C}$ is an ideal of morphisms that is given by subgroups
$\operatorname{Rad}(X,Y)\subseteq\operatorname{Hom}(X,Y)$
for each pair of objects $X,Y$ such that $\operatorname{Rad}(X,X)$ equals the Jacobson radical of the endomorphism ring $\operatorname{End}(X)$. A useful consequence of (2.1) is that for any pair of objects $X,Y\in\mathcal{A}$ we have
(2.2) $P(X)\cong P(Y)\qquad\iff\qquad E(X)\cong E(Y)$
where $E(X)$ denotes an injective envelope of $X$; see also [11, Satz 1.5].
The following lemma shows that the spectral category controls the decomposition of injective objects.
Lemma 2.1.
Let $X\in\mathcal{A}$ be an injective object and $P(X)=Y_{1}\oplus Y_{2}$ a decomposition. Then there exists a decomposition $X=X_{1}\oplus X_{2}$ such that $Y_{i}=P(X_{i})$ for $i=1,2$.
Proof.
There is a decomposition into idempotent endomorphisms $\operatorname{id}_{P(X)}=\phi_{1}+\phi_{2}$ such that $Y_{i}=\operatorname{Ker}\phi_{i}$. Using (2.1) we decompose $\operatorname{id}_{X}=\psi_{1}+\psi_{2}$ such that $P(\psi_{i})=\phi_{i}$. Now set $X_{i}=E(\operatorname{Ker}\psi_{i})\subseteq X$ and apply (2.2). ∎
Spectral categories and their structure are closely related to (modules over) self-injective von Neumann regular rings, because the spectral categories are precisely (up to equivalence) the categories of nonsingular injective modules over such rings. This is explained in [11, 30], and for the structure of such rings we refer to [13]. [^1] Given a spectral category and a generator $G$, the endomorphism ring of $G$ is self-injective von Neumann regular, and the equivalence is induced by $\operatorname{Hom}(G,-)$. There is an elaborated decomposition theory for operator algebras which uses certain types and goes back to Murray, von Neumann, and Kaplansky [20, 26]. The analogous decomposition theory for spectral categories is developed in [14, 30]. For some further structure of spectral categories, see [8].
Localising subcategories
Let $\mathcal{A}$ be a Grothendieck category. A full subcategory of $\mathcal{A}$ is localising if it is closed under subobjects, quotients, extensions, and arbitrary coproducts. The localising subcategories of $\mathcal{A}$ form a set, because for a fixed generator $G\in\mathcal{A}$ any localising subcategory $\mathcal{U}\subseteq\mathcal{A}$ is determined by the set of quotients $G/G^{\prime}$ which belong to $\mathcal{U}$. The localising subcategories are partially ordered by inclusion and closed under arbitrary intersections; so they form a complete lattice which we denote by $\mathbf{L}(\mathcal{A})$.
For a localising subcategory $\mathcal{U}\subseteq\mathcal{A}$ and an object $X\in\mathcal{A}$ let $t_{\mathcal{U}}(X)\subseteq X$ denote the maximal subobject in $\mathcal{U}$. The following lemma provides a way of describing the join in $\mathbf{L}(\mathcal{A})$.
Lemma 2.2.
Let $(\mathcal{U}_{\alpha})$ be a family in $\mathbf{L}(\mathcal{A})$ and set $\mathcal{U}=\bigvee_{\alpha}\mathcal{U}_{\alpha}$. For any $X\in\mathcal{A}$ we have
$\sum_{\alpha}t_{\mathcal{U}_{\alpha}}(X)=t_{\mathcal{U}}(X).$
In particular,
$t_{\mathcal{U}}(X)=0\qquad\iff\qquad t_{\mathcal{U}_{\alpha}}(X)=0\quad\text{for all}\quad\alpha.$
Proof.
We begin with the second assertion. Let $X\to E(X)$ denote an injective envelope. Then $t_{\mathcal{U}}(X)=0$ if and only if $t_{\mathcal{U}}(E(X))=0$. So we may assume that $X$ is injective. But then the assertion is clear, since $t_{\mathcal{U}}(X)=0$ if and only if $\operatorname{Hom}(U,X)=0$ for all $U\in\mathcal{U}$, and the kernel of $\operatorname{Hom}(-,X)$ is a localising subcategory.
Now consider the subobjects
$X^{\prime}:=\sum_{\alpha}t_{\mathcal{U}_{\alpha}}(X)\subseteq t_{\mathcal{U}}(X)\subseteq X$
If $X$ is in $\mathcal{U}$, then $t_{\mathcal{U}_{\alpha}}(X/X^{\prime})=0$ for all $\alpha$, and therefore the first part of the proof yields $t_{\mathcal{U}}(X/X^{\prime})=0$. Thus $X^{\prime}=X$. ∎
A complete lattice is a frame if finite meets distribute over arbitrary joins, and it is a coframe if finite joins distribute over arbitrary meets. We record the following well known fact; see for instance [12] for the case of a module category, and we refer to [34] for a proof that uses an explicit construction of the join in the lattice of localising subcategories.
Proposition 2.3.
Let $\mathcal{A}$ be a Grothendieck category. For an element $\mathcal{U}$ and a family $(\mathcal{V}_{\alpha})$ in $\mathbf{L}(\mathcal{A})$ we have
$\mathcal{U}\wedge\left(\bigvee_{\alpha}\mathcal{V}_{\alpha}\right)=\bigvee_{\alpha}(\mathcal{U}\wedge\mathcal{V}_{\alpha}).$
Thus the lattice $\mathbf{L}(\mathcal{A})$ is a frame.
Proof.
The inclusion $\supseteq$ is automatic; so we need to show $\subseteq$. Let $X$ be an object in $\mathcal{U}\wedge\left(\bigvee_{\alpha}\mathcal{V}_{\alpha}\right)$. Then Lemma 2.2 implies $X=\sum_{\alpha}t_{\mathcal{V}_{\alpha}}(X)$, and therefore $X$ belongs to $\bigvee_{\alpha}(\mathcal{U}\wedge\mathcal{V}_{\alpha})$ since $t_{\mathcal{V}_{\alpha}}(X)\in\mathcal{U}\wedge\mathcal{V}_{\alpha}$ for each $\alpha$. ∎
Proposition 2.4.
Let $\mathcal{A}$ be a spectral Grothendieck category. For an element $\mathcal{U}$ and a family $(\mathcal{V}_{\alpha})$ in $\mathbf{L}(\mathcal{A})$ we have
$\mathcal{U}\vee\left(\bigwedge_{\alpha}\mathcal{V}_{\alpha}\right)=\bigwedge_{\alpha}(\mathcal{U}\vee\mathcal{V}_{\alpha}).$
Thus the lattice $\mathbf{L}(\mathcal{A})$ is a coframe.
Proof.
The inclusion $\subseteq$ is automatic; so we need to show $\supseteq$. Any object $X$ can be written as $X=t_{\mathcal{U}}(X)\oplus X/t_{\mathcal{U}}(X)$. A simple calculation shows that $X\in\mathcal{U}\vee\mathcal{V}_{\alpha}$ if and only if $X/t_{\mathcal{U}}(X)\in\mathcal{V}_{\alpha}$. Thus $X\in\bigwedge_{\alpha}(\mathcal{U}\vee\mathcal{V}_{\alpha})$ implies $X/t_{\mathcal{U}}(X)\in\bigwedge_{\alpha}\mathcal{V}_{\alpha}$, and therefore $X\in\mathcal{U}\vee\left(\bigwedge_{\alpha}\mathcal{V}_{\alpha}\right)$. ∎
Discrete versus continuous
Let $\mathcal{A}$ be a spectral category. The category $\mathcal{A}$ is called discrete if every object is a coproduct of simple objects. Any object $X\in\mathcal{A}$ admits a decomposition
$X=\operatorname{soc}(X)\oplus\operatorname{rad}(X),$
where the socle $\operatorname{soc}(X)$ is the sum of all simple subobjects and the radical $\operatorname{rad}(X)$ is the intersection of all maximal subobjects. We set
$\mathcal{A}_{\mathrm{d}}:=\{X\in\mathcal{A}\mid\operatorname{soc}(X)=X\}\qquad\text{and}\qquad\mathcal{A}_{\mathrm{c}}:=\{X\in\mathcal{A}\mid\operatorname{rad}(X)=X\}.$
This yields a decomposition $\mathcal{A}=\mathcal{A}_{\mathrm{d}}\times\mathcal{A}_{\mathrm{c}}$ of spectral Grothendieck categories such that $\mathcal{A}_{\mathrm{d}}$ is discrete and $(\mathcal{A}_{\mathrm{c}})_{\mathrm{d}}=0=(\mathcal{A}_{\mathrm{d}})_{\mathrm{c}}$; see [11, §3].
Lemma 2.5.
Let $\mathcal{A}$ be spectral. The assignment $\mathcal{U}\mapsto\mathcal{U}\cap\operatorname{Sp}\mathcal{A}$ induces a lattice isomorphism between $\mathbf{L}(\mathcal{A}_{\mathrm{d}})$ and the lattice of subsets of $\operatorname{Sp}\mathcal{A}$.
Proof.
The inverse map sends a subset $\mathcal{V}\subseteq\operatorname{Sp}\mathcal{A}$ to the subcategory of $\mathcal{A}$ consisting of all coproducts of objects in $\mathcal{V}$. ∎
Example 2.6.
Let $\mathcal{A}$ be a locally noetherian Grothendieck category. Then its spectral category $\operatorname{Spec}\mathcal{A}$ is discrete [10, IV.2]. Thus in this case $\mathbf{L}(\operatorname{Spec}\mathcal{A})$ identifies with the lattice of subsets of $\operatorname{Sp}\mathcal{A}$.
The Boolean spectrum
Let $\mathcal{A}$ be a Grothendieck category. For a subcategory $\mathcal{C}\subseteq\mathcal{A}$ we set
$\mathcal{C}^{\perp}$ :=$:=\{X\in\mathcal{A}\mid\operatorname{Hom}(C,X)=0\text{ for all }C\in\mathcal{C}\},$
${}^{\perp}\mathcal{C}$ :=$:=\{X\in\mathcal{A}\mid\operatorname{Hom}(X,C)=0\text{ for all }C\in\mathcal{C}\}.$
Localising subcategories of $\mathcal{A}$ are in bijective correspondence to hereditary torsion pairs, by taking a subcategory $\mathcal{C}$ to the pair $(\mathcal{C},\mathcal{C}^{\perp})$; see [33, VI.3]. Torsion pairs for spectral categories admit the following elementary description.
Lemma 2.7.
Let $\mathcal{A}$ be a spectral Grothendieck category. A pair $(\mathcal{C},\mathcal{D})$ of full subcategories is a torsion pair if and only if
  1. (1)
    $\mathcal{C}$ and $\mathcal{D}$ are closed under direct summands,
  2. (2)
    $\mathcal{C}\cap\mathcal{D}=0$, and
  3. (3)
    $\mathcal{A}=\{C\oplus D\mid C\in\mathcal{C},\,D\in\mathcal{D}\}$.
Proof.
Clearly, for any torsion pair $(\mathcal{C},\mathcal{D})$ the conditions (1)–(3) hold. Conversely, these conditions imply that $\mathcal{C}^{\perp}=\mathcal{D}$ and $\mathcal{C}={{}^{\perp}\mathcal{D}}$. Thus $(\mathcal{C},\mathcal{D})$ is a torsion pair. ∎
We record the above in lattice-theoretical terms. Recall that a bounded lattice $L$ is complemented if for each $x\in L$ there exists a complement $y\in L$, so $x\wedge y=0$ and $x\vee y=1$. Such a complement is unique when $L$ is distributive, and then we write $\neg x$ for the complement of $x$. A complemented distributive lattice is called Boolean lattice.
Proposition 2.8.
Let $\mathcal{A}$ be a spectral Grothendieck category. Then the lattice $\mathbf{L}(\mathcal{A})$ is a complete Boolean lattice.
Proof.
First observe that all torsion pairs are hereditary since $\mathcal{A}$ is spectral. Now let $(\mathcal{C},\mathcal{D})$ be a torsion pair. Then it follows from Lemma 2.7 that $(\mathcal{D},\mathcal{C})$ is again a torsion pair. Thus $\mathcal{D}=\mathcal{C}^{\perp}$ is a complement for $\mathcal{C}\in\mathbf{L}(\mathcal{A})$. ∎
Let us collect some basic facts from Stone duality; for details and proofs see [7, 18]. Let $A=(A,\vee,\wedge)$ be a Boolean lattice. We may identify $A$ with the corresponding Boolean ring $(A,+,\cdot)$, with operations given by
$x+y:=(x\wedge\neg y)\vee(y\wedge\neg x)\qquad\text{and}\qquad x\cdot y:=x\wedge y,$
and keeping in mind that the category of Boolean lattices is isomorphic to the category of Boolean rings. Let $\operatorname{Spec}(A)$ denote the spectrum of prime ideals of $A$ with the Zariski topology. For a topological space $T$ we write $\operatorname{Clop}(T)$ for its Boolean lattice of clopen subsets. This yields an adjoint pair of contravariant functors
${\mathbf{Top}}$${\mathbf{BRing}}$$\scriptstyle{\operatorname{Clop}=\operatorname{Hom}(-,\mathbf{2})}$$\scriptstyle{\operatorname{Spec}=\operatorname{Hom}(-,\mathbf{2})}$
where $\mathbf{2}=\{0,1\}$ is either viewed as topological space with the discrete topology, or $\mathbf{2}=\mathbb{F}_{2}$ is viewed as a Boolean ring. This means there is a natural bijection
(2.3) $\operatorname{Hom}(A,\operatorname{Clop}(T))\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Hom}(T,\operatorname{Spec}(A))$
which sends $\phi\colon A\to\operatorname{Clop}(T)$ to the map $T\to\operatorname{Spec}(A)$ given by
$p\longmapsto\{x\in A\mid p\not\in\phi(x)\}\qquad\text{for }p\in T.$
Stone duality implies that for any Boolean ring $A$ the unit of the adjunction
(2.4) $A\longrightarrow\operatorname{Clop}(\operatorname{Spec}(A)),\quad x\mapsto\{\mathfrak{p}\in\operatorname{Spec}(A)\mid x\not\in\mathfrak{p}\}$
is an isomorphism.
Definition 2.9.
The Boolean spectrum of a Grothendieck category $\mathcal{A}$ is the space
$\operatorname{Spc}\mathcal{A}:=\operatorname{Spec}(\mathbf{L}(\operatorname{Spec}\mathcal{A})).$
For a Grothendieck category $\mathcal{A}$ the decomposition $\operatorname{Spec}\mathcal{A}=(\operatorname{Spec}\mathcal{A})_{\mathrm{d}}\times(\operatorname{Spec}\mathcal{A})_{\mathrm{c}}$ implies that $\operatorname{Spc}\mathcal{A}$ equals the disjoint union
(2.5) $\operatorname{Spc}\mathcal{A}=\operatorname{Spec}(\mathbf{L}((\operatorname{Spec}\mathcal{A})_{\mathrm{d}}))\sqcup\operatorname{Spec}(\mathbf{L}((\operatorname{Spec}\mathcal{A})_{\mathrm{c}}))$
where
$\operatorname{Spec}(\mathbf{L}((\operatorname{Spec}\mathcal{A})_{\mathrm{d}}))\cong\operatorname{Sp}\mathcal{A}$
carries the discrete topology. In particular, we see that the set $\operatorname{Sp}\mathcal{A}$ of isoclasses of indecomposable injective objects embeds canonically into $\operatorname{Spc}\mathcal{A}$; cf. Lemma 2.5.
Remark 2.10.
It would be interesting to know which Boolean lattices arise from spectral categories. More specifically, given a Boolean lattice $L$, is there a procedure to construct a self-injective von Neumann regular ring $\Lambda$ such that for the category $\mathcal{A}$ of singular injective $\Lambda$-modules we have $\mathbf{L}(\mathcal{A})\cong L$? There is an elaborated structure theory for self-injective von Neumann regular rings, which seems to neglect procedures for actual constructions.
Coproduct decompositions
We provide an explicit description of the join in the lattice $\mathbf{L}(\mathcal{A})$ when $\mathcal{A}$ is spectral. We use the following property of a Boolean lattice.
Lemma 2.11.
Let $A$ be a complete Boolean lattice and $x=\bigvee_{\alpha}x_{\alpha}$ an element. Then there are elements $y_{\alpha}\leq x_{\alpha}$ such that $x=\bigvee_{\alpha}y_{\alpha}$ and $y_{\alpha}\wedge y_{\beta}=0$ for all $\alpha\neq\beta$.
Proof.
We consider the set of families $(y_{\alpha})$ of objects in $A$ such that $y_{\alpha}\leq x_{\alpha}$ and $y_{\alpha}\wedge y_{\beta}=0$ for all $\alpha\neq\beta$. This set is partially ordered via $(y_{\alpha})\leq(y^{\prime}_{\alpha})$ if $y_{\alpha}\leq y^{\prime}_{\alpha}$ for all $\alpha$. An application of Zorn’s lemma yields a maximal element and it is easily checked that $x=\bigvee_{\alpha}y_{\alpha}$ for such maximal $(y_{\alpha})$. ∎
Proposition 2.12.
Let $\mathcal{A}$ be a spectral Grothendieck category. For a family $(\mathcal{U}_{\alpha})$ in $\mathbf{L}(\mathcal{A})$ we have
$\bigvee_{\alpha}\mathcal{U}_{\alpha}=\left\{\coprod_{\alpha}X_{\alpha}\mid X_{\alpha}\in\mathcal{U}_{\alpha}\text{ for all }\alpha\right\}.$
Proof.
Set $\mathcal{U}=\bigvee_{\alpha}\mathcal{U}_{\alpha}$. We appy Lemma 2.11 and obtain a decomposition $\mathcal{U}=\bigvee_{\alpha}\mathcal{V}_{\alpha}$ such that $\mathcal{V}_{\alpha}\subseteq\mathcal{U}_{\alpha}$ and $\mathcal{V}_{\alpha}\cap\mathcal{V}_{\beta}=0$ for all $\alpha\neq\beta$. Given $X\in\mathcal{U}$ we have $X=\sum_{\alpha}t_{\mathcal{V}_{\alpha}}(X)$ by Lemma 2.2. Moreover, $t_{\mathcal{V}_{\beta}}(X)\cap\sum_{\alpha\neq\beta}t_{\mathcal{V}_{\alpha}}(X)=0$ for all $\beta$ since $\mathcal{V}_{\beta}\wedge\left(\bigvee_{\alpha\neq\beta}\mathcal{V}_{\alpha}\right)=0$. Thus $X=\coprod_{\alpha}t_{\mathcal{V}_{\alpha}}(X)$. ∎
Functoriality
We briefly discuss the functoriality of the assignment $\mathcal{A}\mapsto\operatorname{Spec}\mathcal{A}$. An exact and coproduct preserving functor $\mathcal{A}\to\mathcal{B}$ between Grothendieck categories admits a right adjoint which restricts to an additive functor $\operatorname{Inj}\mathcal{B}\to\operatorname{Inj}\mathcal{A}$. Using (2.1) one obtains an exact functor $\operatorname{Spec}\mathcal{B}\to\operatorname{Spec}\mathcal{A}$.
Now let $\mathcal{A}^{\prime}\subseteq\mathcal{A}$ be a localising subcategory and set $\mathcal{A}^{\prime\prime}=\mathcal{A}/\mathcal{A}^{\prime}$.
Lemma 2.13.
The exact sequence $\mathcal{A}^{\prime}\rightarrowtail\mathcal{A}\twoheadrightarrow\mathcal{A}^{\prime\prime}$ induces an exact sequence
$\operatorname{Spec}\mathcal{A}^{\prime\prime}\rightarrowtail\operatorname{Spec}\mathcal{A}\twoheadrightarrow\operatorname{Spec}\mathcal{A}^{\prime}.$
These functors admit right adjoints and yield decompositions
$\operatorname{Spec}\mathcal{A}=\operatorname{Spec}\mathcal{A}^{\prime}\times\operatorname{Spec}\mathcal{A}^{\prime\prime}\qquad\text{and}\qquad\operatorname{Spc}\mathcal{A}=\operatorname{Spc}\mathcal{A}^{\prime}\sqcup\operatorname{Spc}\mathcal{A}^{\prime\prime}.$
Proof.
The functor $\operatorname{Spec}\mathcal{A}\to\operatorname{Spec}\mathcal{A}^{\prime}$ is induced by $t_{\mathcal{A}^{\prime}}\colon\mathcal{A}\to\mathcal{A}^{\prime}$. It is easily checked that $\operatorname{Spec}\mathcal{A}^{\prime\prime}\to\operatorname{Spec}\mathcal{A}$ identifies $\operatorname{Spec}\mathcal{A}^{\prime\prime}$ with the kernel of $\operatorname{Spec}\mathcal{A}\to\operatorname{Spec}\mathcal{A}^{\prime}$, which is closed under arbitrary coproducts. The inclusion $\mathcal{A}^{\prime}\to\mathcal{A}$ preserves essential extensions and induces therefore a functor $\operatorname{Spec}\mathcal{A}^{\prime}\to\operatorname{Spec}\mathcal{A}$ which is a right inverse for $\operatorname{Spec}\mathcal{A}\to\operatorname{Spec}\mathcal{A}^{\prime}$. Thus $\operatorname{Spec}\mathcal{A}\to\operatorname{Spec}\mathcal{A}^{\prime}$ is a localisation functor, and this yields the decomposition $\operatorname{Spec}\mathcal{A}=\operatorname{Spec}\mathcal{A}^{\prime}\times\operatorname{Spec}\mathcal{A}^{\prime\prime}$. The functor from spectral categories to spaces given by $\mathcal{C}\mapsto\operatorname{Spec}\mathbf{L}(\mathcal{C})$ is contravariant and takes products to coproducts. ∎
Monoidal structure
Let $\mathcal{A}$ be a spectral Grothendieck category. Any choice of a generator $G\in\mathcal{A}$ induces a tensor product $-\otimes_{G}-\colon\mathcal{A}\times\mathcal{A}\to\mathcal{A}$ as follows. We write $\mathcal{G}$ for the full subcategory given by the direct summands of finite direct sums of copies of $G$. Let $-\otimes_{G}-\colon\mathcal{G}\times\mathcal{G}\to\mathcal{G}$ be the unique additive functor such that $C\otimes_{G}G=C=G\otimes_{G}C$ for all $C\in\mathcal{G}$. An arbitrary object $X\in\mathcal{A}$ can be written canonically as (filtered) colimit $X=\operatorname*{colim}_{C\to X}C$ where $C\to X$ runs through all morphisms from objects in $\mathcal{G}$. Then set
$X\otimes_{G}Y:=\operatorname*{colim}_{\begin{subarray}{c}C\to X\\ D\to Y\end{subarray}}C\otimes_{G}D.$
It is clear that the $X\otimes_{G}-$ is exact and preserves all coproducts. Also, any localising subcategory is a tensor ideal. Let $\langle X\rangle$ denote the localising subcategory generated by $X$.
Lemma 2.14.
For a pair of objects $X,Y$ in $\mathcal{A}$ we have $\langle X\rangle\cap\langle X\rangle=\langle X\otimes_{G}Y\rangle$.
Proof.
We have $t_{\langle X\rangle}=t_{\langle X\rangle}(G)\otimes_{G}-$. Thus for an object $M\in\langle X\rangle\cap\langle X\rangle$ we have
$M\cong t_{\langle X\rangle}(G)\otimes_{G}t_{\langle Y\rangle}(G)\otimes_{G}M\in\langle t_{\langle X\rangle}(G)\otimes_{G}t_{\langle Y\rangle}(G)\rangle\subseteq\langle X\otimes_{G}Y\rangle.$
The other inclusion is clear. ∎
3. Essentially closed subcategories
We fix a Grothendieck category $\mathcal{A}$ and introduce a class of subcategories which is controlled by its spectral category. This yields a notion of support for the objects of $\mathcal{A}$. To simplify notation we set
$\mathbf{S}(\mathcal{A}):=\mathbf{L}(\operatorname{Spec}\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Clop}(\operatorname{Spc}\mathcal{A}).$
Essentially closed subcategories and support
We begin with the main definition of this section; the terminology is justified by Theorem 3.3, because we look at the operations that are preserved by the functor $\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ inverting all essential monomorphisms.
Definition 3.1.
A full subcategory of $\mathcal{A}$ is called essentially closed if it is closed under arbitrary coproducts, subobjects, and essential extensions.
The essentially closed subcategories are partially ordered by inclusion and closed under arbitrary intersections; so they form a complete lattice. When $\mathcal{A}$ is spectral, then its essentially closed subcategories are precisely its localising subcategories. For a class $\mathcal{X}\subseteq\mathcal{A}$ let $\langle\mathcal{X}\rangle$ denote the smallest essentially closed subcategory of $\mathcal{A}$ that contains $\mathcal{X}$.
Lemma 3.2.
Let $\mathcal{A}$ be a Grothendieck category.
  1. (1)
    If $\mathcal{U}\subseteq\operatorname{Spec}\mathcal{A}$ is localising, then $P^{-1}(\mathcal{U})\subseteq\mathcal{A}$ is essentially closed.
  2. (2)
    If $\mathcal{C}\subseteq\mathcal{A}$ is essentially closed, then $P(\mathcal{C})\subseteq\operatorname{Spec}\mathcal{A}$ is localising.
Proof.
(1) Let $X\subseteq Y$ be a subobject in $\mathcal{A}$ with $P(Y)\in\mathcal{U}$. Then $P(X)\in\mathcal{U}$ since $P$ is left exact. Thus $P^{-1}(\mathcal{U})$ is closed under subobjects. Analogously, $P^{-1}(\mathcal{U})$ is closed under coproducts since $P$ preserves coproducts, and it is closed under essential extensions as $P$ maps those to isomorphisms.
(2) First observe that the subcategory $P(\mathcal{C})$ is replete, so given a pair of objects $X,Y\in\mathcal{A}$ such that $P(X)\cong P(Y)$, then $X\in\mathcal{C}$ implies $Y\in\mathcal{C}$. This follows from (2.2), because when $X$ and $Y$ have isomorphic injective envelopes, then $X\in\mathcal{C}$ implies $Y\in\mathcal{C}$. Next we show that $P(\mathcal{C})$ is closed under coproducts and subobjects. Clearly, $P(\mathcal{C})$ is closed under coproducts since $P$ preserves coproducts. Any subobject $Y\subseteq P(X)$ of an object in $P(\mathcal{C})$ is actually a direct summand of $P(X)$, and we may assume that $X$ is injective. Using the equivalence (2.1), we have $Y\cong\operatorname{Ker}P(\phi)\cong P(\operatorname{Ker}\phi)$ for some morphism $\phi\colon X\to X$. Thus $P(\mathcal{C})$ is closed under subobjects. ∎
For an object $X\in\mathcal{A}$ its support is by definition
(3.1) $\operatorname{supp}(X):=\inf\{\mathcal{U}\in\mathbf{S}(\mathcal{A})\mid P(X)\in\mathcal{U}\}=\langle P(X)\rangle.$
For a class $\mathcal{X}\subseteq\mathcal{A}$ we set $\operatorname{supp}(\mathcal{X}):=\bigvee_{X\in\mathcal{X}}\operatorname{supp}(X)$ and note that
(3.2) $\operatorname{supp}(\mathcal{X})=\bigvee_{X\in\mathcal{X}}\langle P(X)\rangle=\langle P(\mathcal{X})\rangle.$
Theorem 3.3.
Let $\mathcal{A}$ be a Grothendieck category. The assignment
$\mathcal{C}\longmapsto\operatorname{supp}(\mathcal{C})=P(\mathcal{C})$
induces a lattice isomorphism
$\{\mathcal{C}\subseteq\mathcal{A}\mid\mathcal{C}\text{ essentially closed}\}\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{S}(\mathcal{A}).$
The inverse map is given by $\mathcal{U}\mapsto P^{-1}(\mathcal{U})$.
Proof.
The functor $P$ equals the identity on objects. From this and Lemma 3.2 it follows that $\mathcal{C}\mapsto P(\mathcal{C})$ and $\mathcal{U}\mapsto P^{-1}(\mathcal{U})$ provide mutually inverse bijections between the essentially closed subcategories of $\mathcal{A}$ and $\operatorname{Spec}\mathcal{A}$, respectively. The equality $\operatorname{supp}(\mathcal{C})=P(\mathcal{C})$ follows from (3.2). ∎
In this generality the theorem seems to be new. Special cases were studied before; see the discussion in Remark 4.7 when we establish another correspondence for subcategories of $\mathcal{A}$. Essentially closed subcategories were introduced for module categories by Dauns under the name ‘saturated class’ [6]. He showed that they form a Boolean lattice, though his proof is different from ours and does not use the spectral category. Saturated classes were used in [6] to develop a decomposition theory for modules, analogous to the one for von Neumann regular rings. Proposition 3.6 gives some flavour of such decomposition results.
Corollary 3.4.
Every essentially closed subcategory of $\mathcal{A}$ is of the form $\langle X\rangle$ for some object $X\in\mathcal{A}$.
Proof.
The assertion is clear when $\mathcal{A}$ is spectral, because any localising subcategory is a Grothendieck category and contains therefore a generator. When $\mathcal{A}$ is arbitrary and $\mathcal{C}\subseteq\mathcal{A}$ is essentially closed, we can choose $X\in\mathcal{A}$ with $P(\mathcal{C})=\langle P(X)\rangle$. This implies $\mathcal{C}=\langle X\rangle$. ∎
Remark 3.5.
The complement of an essentially closed subcategory $\mathcal{C}\subseteq\mathcal{A}$ admits an easy description: it is the essentially closed subcategory consisting of all objects $X\in\mathcal{A}$ such that any subobject of $X$ is zero when it belongs to $\mathcal{C}$.
Coproduct decompositions
A coproduct decomposition of an object in $\mathcal{A}$ induces via the assignment $X\mapsto\langle X\rangle$ a decomposition in the lattice of essentially closed subcategories. This correspondence works in both directions as follows.
Proposition 3.6.
Let $\mathcal{A}$ be a Grothendieck category and $X\in\mathcal{A}$ an object.
  1. (1)
    A decomposition $X=\coprod_{\alpha}X_{\alpha}$ in $\mathcal{A}$ induces a decomposition $\langle X\rangle=\bigvee_{\alpha}\langle X_{\alpha}\rangle$ into essentially closed subcategories.
  2. (2)
    Suppose that $X$ is injective and let $\langle X\rangle=\bigvee_{\alpha}\mathcal{C}_{\alpha}$ be a decomposition into essentially closed subcategories. Then there are direct summands $X_{\alpha}\subseteq X$ such that $X=E(\coprod_{\alpha}X_{\alpha})$ and $X_{\alpha}\in\mathcal{C}_{\alpha}$ for all $\alpha$. Moreover, $\langle X_{\alpha}\rangle=\mathcal{C}_{\alpha}$ for all $\alpha$ provided that $\mathcal{C}_{\alpha}\wedge\mathcal{C}_{\beta}=0$ for all $\alpha\neq\beta$.
Proof.
(1) is clear from the definitions. For (2) we may assume that $\mathcal{C}_{\alpha}\wedge\mathcal{C}_{\beta}=0$ for all $\alpha\neq\beta$ by replacing $\mathcal{C}_{\alpha}$ with a smaller subcategory if needed; see Lemma 2.11. Suppose first that $\mathcal{A}$ is spectral. Then there is an essentially unique decomposition $X=\coprod_{\alpha}X_{\alpha}$ in $\mathcal{A}$ with $\langle X_{\alpha}\rangle=\mathcal{C}_{\alpha}$ for all $\alpha$, by choosing $X_{\alpha}=t_{\mathcal{C}_{\alpha}}(X)$. This follows from Lemma 2.2; see also Proposition 2.12. When $\mathcal{A}$ is arbitrary and $X$ is injective we obtain a decomposition of $X$ by decomposing $P(X)$ in $\operatorname{Spec}\mathcal{A}$ and applying Theorem 3.3. More precisely, for each $\alpha$ we apply Lemma 2.1 and choose a direct summand $X_{\alpha}\subseteq X$ such that $P(X_{\alpha})=t_{P(\mathcal{C}_{\alpha})}(P(X))$. Then the equality $X=E(\coprod_{\alpha}X_{\alpha})$ follows from (2.2). ∎
Remark 3.7.
The identity $X=E(\coprod_{\alpha}X_{\alpha})$ in Proposition 3.6 simplifies to $X=\coprod_{\alpha}X_{\alpha}$ when the index set is finite or when $\mathcal{A}$ is locally noetherian.
Support data
The notion of a support datum for tensor triangulated categories was introduced by Balmer in [1]. The following is the analogue for Grothendieck categories, where the intersection of essentially closed subcategories plays the role of the tensor product. [^2] This is reasonable since $\langle X\otimes Y\rangle=\langle X\rangle\cap\langle Y\rangle$ for objects $X,Y$ in a tensor triangulated category, where $\langle X\rangle$ denotes the radical thick tensor ideal generated by $X$. See also Lemma 2.14 for the case of a spectral category.
Definition 3.8.
A support datum on a Grothendieck category $\mathcal{A}$ is a pair $(T,\sigma)$ consisting of a topological space $T$ and a map $\sigma\colon\operatorname{Ob}\mathcal{A}\to\operatorname{Clop}(T)$ such that
  1. (S00)
    $\sigma(X)=\varnothing$ for $X=0$,
  2. (S1111)
    $\sigma(X)=T$ for every generator $X$ of $\mathcal{A}$,
[…]
    $\sigma(X)\cup\sigma(Y)=\sigma(X\oplus Y)$ for all objects $X,Y$ in $\mathcal{A}$, and
[…]
    $\sigma(X)\cap\sigma(Y)=\sigma(Z)$ for all objects $X,Y,Z$ in $\mathcal{A}$ with $\langle X\rangle\cap\langle Y\rangle=\langle Z\rangle$.
A morphism $(T,\sigma)\to(T^{\prime},\sigma^{\prime})$ is given by a continuous map $f\colon T\to T^{\prime}$ such that $\sigma(X)=f^{-1}(\sigma^{\prime}(X))$ for all $X\in\mathcal{A}$.
It is convenient to reformulate the notion of a support datum as follows.
Lemma 3.9.
Let $(T,\sigma)$ be a support datum. Given objects $X,Y$ in $\mathcal{A}$, the condition $\langle X\rangle\subseteq\langle Y\rangle$ implies $\sigma(X)\subseteq\sigma(Y)$. Therefore the assignment $\langle X\rangle\mapsto\sigma(X)$ induces a map
$\{\mathcal{C}\subseteq\mathcal{A}\mid\mathcal{C}\text{ essentially closed}\}\longrightarrow\operatorname{Clop}(T)$
which is a lattice homomorphism.
Proof.
[…]
By slight abuse of notation and using the isomorphism (2.4) we write $\operatorname{supp}$ for the composite
$\operatorname{Ob}\mathcal{A}\xrightarrow{\ \operatorname{supp}\ }\mathbf{S}(\mathcal{A})\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\operatorname{Clop}(\operatorname{Spec}(\mathbf{S}(\mathcal{A})))=\operatorname{Clop}(\operatorname{Spc}\mathcal{A}).$
Theorem 3.10.
The pair $(\operatorname{Spc}\mathcal{A},\operatorname{supp})$ is a support datum on $\mathcal{A}$. Moreover, for any support datum $(T,\sigma)$ there is a unique continuous map $f\colon T\to\operatorname{Spc}\mathcal{A}$ such that $\sigma(X)=f^{-1}(\operatorname{supp}(X))$ for all $X\in\mathcal{A}$. Explicitly, the map $f$ is given by the adjunction (2.3) and therefore
$f(p)=\{\langle X\rangle\in\mathbf{S}(\mathcal{A})\mid p\not\in\sigma(X)\}\qquad\text{for }p\in T.$
Proof.
It is clear that $(\operatorname{Spc}\mathcal{A},\operatorname{supp})$ is a support datum. Each support datum $(T,\sigma)$ induces a homomorphism $\{\mathcal{C}\subseteq\mathcal{A}\mid\mathcal{C}\text{ essentially closed}\}\to\operatorname{Clop}(T)$ by Lemma 3.9. The composite with the isomorphism $\mathbf{S}(\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\{\mathcal{C}\subseteq\mathcal{A}\mid\mathcal{C}\text{ essentially closed}\}$ from Theorem 3.3 yields a lattice homomorphism, which corresponds via the adjunction (2.3) to a continuous map $T\to\operatorname{Spec}(\mathbf{S}(\mathcal{A}))=\operatorname{Spc}\mathcal{A}$ with the above explicit description. ∎
The clopen subsets of $\operatorname{Spc}\mathcal{A}$ form a complete lattice. This reflects the existence of arbitrary coproducts in $\mathcal{A}$, but it need not hold for an arbitrary space.
Proposition 3.11.
Let $(X_{\alpha})$ be a family of objects in $\mathcal{A}$. Then we have
$\operatorname{supp}\left(\coprod_{\alpha}X_{\alpha}\right)=\bigvee_{\alpha}\operatorname{supp}(X_{\alpha}).$
Proof.
The assertion follows from the equality $\langle\coprod_{\alpha}X_{\alpha}\rangle=\bigvee_{\alpha}\langle X_{\alpha}\rangle$. ∎
We end our discussion of support with a simple observation which demonstrates the benefit of the extension $\operatorname{Sp}\mathcal{A}\hookrightarrow\operatorname{Spc}\mathcal{A}$.
Remark 3.12.
For $X\in\mathcal{A}$ we have $\operatorname{supp}(X)\neq\varnothing$ when $X\neq 0$. This property cannot be expected for a map $\operatorname{Ob}\mathcal{A}\to\operatorname{Sp}\mathcal{A}$ since $\operatorname{Sp}\mathcal{A}$ may be empty.
Localising subcategories
Next we show that the functor $\mathcal{A}\to\operatorname{Spec}\mathcal{A}$ induces an embedding for the frame of localising subcategories. To this end we consider essentially closed subcategories that are closed under arbitrary products because they arise naturally from localising subcategories. The following lemma explains a tight connection.
Lemma 3.13.
For a full subcategory $\mathcal{D}\subseteq\mathcal{A}$ the following are equivalent.
  1. (1)
    The subcategory $\mathcal{D}$ is essentially closed and closed under arbitrary products.
  2. (2)
    There is a localising subcategory $\mathcal{C}\subseteq\mathcal{A}$ such that $\mathcal{C}^{\perp}=\mathcal{D}$.
Proof.
(1) $\Rightarrow$ (2): We set $\mathcal{C}:={{}^{\perp}\mathcal{D}}$ and this is localising since $\mathcal{D}$ is closed under injective envelopes, so $\mathcal{C}={{}^{\perp}(\mathcal{D}\cap\operatorname{Inj}\mathcal{A})}$. It remains to note that $\mathcal{C}^{\perp}=\mathcal{D}$ where one uses that $\mathcal{D}$ is closed under arbitrary products; see [33, Proposition VI.2.2].
(2) $\Rightarrow$ (1): This is clear. ∎
We continue with a sequence of technical lemmas.
Lemma 3.14.
Let $(\mathcal{C}_{\alpha})$ be a finite family of essentially closed subcategories and $X\in\mathcal{A}$ an injective object. Then $X$ belongs to $\bigvee_{\alpha}\mathcal{C}_{\alpha}$ if and only if there is a decomposition $X=\coprod_{\alpha}X_{\alpha}$ such that $X_{\alpha}\in\mathcal{C}_{\alpha}$ for all $\alpha$.
Proof.
This is an immediate consequence of Proposition 3.6. ∎
Lemma 3.15.
Let $(\mathcal{C}_{\alpha})$ be a family of essentially closed subcategories and suppose that each $\mathcal{C}_{\alpha}$ is closed under arbitrary products. Then $\bigwedge_{\alpha}\mathcal{C}_{\alpha}$ is closed under arbitrary products, and $\bigvee_{\alpha}\mathcal{C}_{\alpha}$ is closed under arbitrary products provided the family is finite.
Proof.
For the lattice of essentially closed subcategories the meet is given by taking intersections, and it is clear that being closed under products is preserved when taking arbitrary intersections. Now suppose the family $(\mathcal{C}_{\alpha})$ is finite. Observe that an essentially closed subcategory $\mathcal{C}\subseteq\mathcal{A}$ is closed under products if and only if $\mathcal{C}\cap\operatorname{Inj}\mathcal{A}$ is closed under products. For $\mathcal{C}=\bigvee_{\alpha}\mathcal{C}_{\alpha}$ we note that $\mathcal{C}\cap\operatorname{Inj}\mathcal{A}$ consists of all objects $\coprod_{\alpha}X_{\alpha}$ with $X_{\alpha}\in\mathcal{C}_{\alpha}\cap\operatorname{Inj}\mathcal{A}$ for all $\alpha$, by Lemma 3.14. For an arbitrary family of objects $(X_{i})$ in $\mathcal{C}\cap\operatorname{Inj}\mathcal{A}$, the product
$\prod_{i}X_{i}=\prod_{i}\left(\coprod_{\alpha}X_{i,\alpha}\right)\cong\coprod_{\alpha}\left(\prod_{i}X_{i,\alpha}\right)$
belongs to $\mathcal{C}$ since $\prod_{i}X_{i,\alpha}$ belongs to $\mathcal{C}_{\alpha}$ for all $\alpha$. ∎
Lemma 3.16.
Let $\mathcal{C}\subseteq\mathcal{A}$ be localising. Then every injective object $X\in\mathcal{A}$ admits a decomposition $X=X^{\prime}\oplus X^{\prime\prime}$ such that $t_{\mathcal{C}}(X)\to X^{\prime}$ is an injective envelope and $t_{\mathcal{C}}(X^{\prime\prime})=0$.
Proof.
The inclusion $t_{\mathcal{C}}(X)\to X$ induces a monomorphism $E(t_{\mathcal{C}}(X))\to X$. One takes for $X^{\prime}$ its image and for $X^{\prime\prime}$ its cokernel. ∎
Lemma 3.17.
Let $\mathcal{C}\subseteq\mathcal{A}$ be localising. Then the essential extensions of objects in $\mathcal{C}$ form an essentially closed subcategory $\bar{\mathcal{C}}\subseteq\mathcal{A}$, and the assignment
$(\mathcal{C},\mathcal{D})\mapsto(P(\bar{\mathcal{C}}),P(\mathcal{D}))$
yields a map between the hereditary torsion pairs for $\mathcal{A}$ and $\operatorname{Spec}\mathcal{A}$, respectively.
Proof.
It is easily checked that the essential extensions of objects in $\mathcal{C}$ form an essentially closed subcategory. Now let $(\mathcal{C},\mathcal{D})$ be a hereditary torsion pair for $\mathcal{A}$. From Lemmas 2.7 and 3.16 it follows that $(P(\bar{\mathcal{C}}),P(\mathcal{D}))$ is a hereditary torsion pair for $\operatorname{Spec}\mathcal{A}$. ∎
Theorem 3.18.
Let $\mathcal{A}$ be a Grothendieck category. There is a pair of maps
${\mathbf{L}(\mathcal{A})}$${\mathbf{L}(\operatorname{Spec}\mathcal{A})}$$\scriptstyle{\sigma}$$\scriptstyle{\tau}$
given by $\sigma(\mathcal{C})=\operatorname{supp}(\mathcal{C})$ and $\tau(\mathcal{U})={{}^{\perp}(P^{-1}(\mathcal{U}^{\perp}))}$ such that $\tau\sigma=\operatorname{id}$. Moreover, $\sigma$ preserves arbitrary joins and finite meets.
Proof.
First observe that $\tau$ is well defined, since $P^{-1}(\mathcal{U}^{\perp})$ is an essentially closed subcategory, and therefore ${{}^{\perp}(P^{-1}(\mathcal{U}^{\perp}))}$ is a localising subcategory of $\mathcal{A}$.
Let $\mathcal{C}\subseteq\mathcal{A}$ be localising and $(\mathcal{C},\mathcal{D})$ the corresponding hereditary torsion pair. Then $(P(\bar{\mathcal{C}}),P(\mathcal{D}))$ is a hereditary torsion pair for $\operatorname{Spec}\mathcal{A}$ by Lemma 3.17, with $\mathcal{U}:=\sigma(\mathcal{C})=P(\bar{\mathcal{C}})$. Thus $\mathcal{U}^{\perp}=P(\mathcal{D})$ and then $P^{-1}(\mathcal{U}^{\perp})=P^{-1}P(\mathcal{D})=\mathcal{D}$. It follows that $\tau(\mathcal{U})=\mathcal{C}$.
The assignment $\mathcal{C}\mapsto\mathcal{C}^{\perp}$ identifies the localising subcategories of $\mathcal{A}$ with the essentially closed subcategories that are closed under products, by Lemma 3.13. This map reverses inclusions and we compose it with the isomorphism from Theorem 3.3. Then it follows from Lemma 3.15 that the image of $\sigma$ is closed under arbitrary joins and finite meets. ∎
I am grateful to Sira Gratz and Greg Stevenson for suggesting the following couple of remarks.
Given a frame $L$, its points are the frame morphisms $L\to\mathbf{2}$, where $\mathbf{2}$ denotes the frame with two elements $\{0\leq 1\}$. The frame $L$ has enough points if for all $x\not\leq y$ in $L$ there exists a point $p\colon L\to\mathbf{2}$ such that $p(x)=1$ and $p(y)=0$.
Remark 3.19.
There are examples showing that the frame $\mathbf{L}(\mathcal{A})$ may have no points [34]. This is in contrast to Proposition 2.8 which shows that the bigger frame $\mathbf{L}(\operatorname{Spec}\mathcal{A})$ has enough points.
Remark 3.20.
[…]
4. Cohomologically stable subcategories
An essentially closed subcategory of a Grothendieck category is always closed under subobjects and extensions, but it need not be closed under quotients. For that reason we consider another class of subcategories which is also controlled by its spectral category. This leads to a notion of exact support for any Grothendieck category.
Exact support
Let $\mathcal{A}$ be an exact category. Thus $\mathcal{A}$ is an additive category together with a distinguished class of short exact sequences. Recall that a full subcategory $\mathcal{C}\subseteq\mathcal{A}$ is thick if $\mathcal{C}$ is closed under direct summands and the two out of three condition holds: any short exact sequence belongs to $\mathcal{C}$ if two of its three terms are in $\mathcal{C}$.
A map $\sigma\colon\operatorname{Ob}\mathcal{A}\to L$ with values in a join-semilattice $L$ is called exact if
  1. (1)
    for all objects $X,Y\in\mathcal{A}$ one has $\sigma(X\oplus Y)=\sigma(X)\vee\sigma(Y)$, and
  2. (2)
    for each exact sequence $0\to X_{1}\to X_{2}\to X_{3}\to 0$ in $\mathcal{A}$ one has
    $\sigma(X_{i})\leq\sigma(X_{j})\vee\sigma(X_{k})\qquad\text{when}\qquad\{i,j,k\}=\{1,2,3\}.$
It is easily checked that $\sigma$ is exact if and only if for each $u\in L$
$\mathcal{A}_{u}:=\{X\in\mathcal{A}\mid\sigma(X)\leq u\}$
is a thick subcategory of $\mathcal{A}$.
Now let $\mathcal{A}$ be a Grothendieck category. We consider again the Boolean lattice
$\mathbf{S}(\mathcal{A}):=\mathbf{L}(\operatorname{Spec}\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Clop}(\operatorname{Spc}\mathcal{A}).$
Also, we consider the derived category $\mathbf{D}(\mathcal{A})$ and the right derived functor
$\mathbf{R}P\colon\mathbf{D}(\mathcal{A})\longrightarrow\mathbf{D}(\operatorname{Spec}\mathcal{A}).$
For an object $X\in\mathcal{A}$ its exact support is by definition
(4.1) $\operatorname{supp}_{\mathrm{ex}}(X):=\inf\{\mathcal{U}\in\mathbf{S}(\mathcal{A})\mid\mathbf{R}P(X)\in\mathbf{D}(\mathcal{U})\}\in\mathbf{S}(\mathcal{A}),$
where we identify
$\mathbf{D}(\mathcal{U})=\{Y\in\mathbf{D}(\operatorname{Spec}\mathcal{A})\mid H^{n}(Y)\in\mathcal{U}\text{ for all }n\in\mathbb{Z}\}.$
For $\mathcal{X}\subseteq\mathcal{A}$ we set $\operatorname{supp}_{\mathrm{ex}}(\mathcal{X}):=\bigvee_{X\in\mathcal{X}}\operatorname{supp}_{\mathrm{ex}}(X)$.
Lemma 4.1.
Let $X\in\mathcal{A}$ be an object and $X\to I$ a minimal injective resolution. Then $\operatorname{supp}_{\mathrm{ex}}(X)$ equals the localising subcategory of $\operatorname{Spec}\mathcal{A}$ consisting of all direct summands of coproducts of objects in $\{P(I^{n})\mid n\in\mathbb{Z}\}$.
Proof.
The functor $P$ sends each differential $I^{n}\to I^{n+1}$ to zero because they are radical morphisms; see [24, Proposition 4.3.18]. Thus $H^{n}(\mathbf{R}P(X))\cong P(I^{n})$ for each $n\in\mathbb{Z}$. ∎
Let $\mathcal{U}\in\mathbf{S}(\mathcal{A})$. We set
$\operatorname{Inj}_{\mathcal{U}}\mathcal{A}:=\{I\in\operatorname{Inj}\mathcal{A}\mid P(I)\in\mathcal{U}\}$
and $\mathcal{A}_{\mathcal{U}}$ denotes the full subcategory of objects $X\in\mathcal{A}$ that admit an injective resolution $X\to I$ with $I^{n}\in\operatorname{Inj}_{\mathcal{U}}\mathcal{A}$ for all $n\in\mathbb{Z}$, so
$\mathcal{A}_{\mathcal{U}}=\{X\in\mathcal{A}\mid\mathbf{R}P(X)\in\mathbf{D}(\mathcal{U})\}=\{X\in\mathcal{A}\mid\operatorname{supp}_{\mathrm{ex}}(X)\subseteq\mathcal{U}\}.$
Lemma 4.2.
For each $\mathcal{U}\in\mathbf{S}(\mathcal{A})$ the subcategory $\mathcal{A}_{\mathcal{U}}$ is thick, and the function
$\operatorname{supp}_{\mathrm{ex}}\colon\operatorname{Ob}\mathcal{A}\longrightarrow\mathbf{S}(\mathcal{A})$
is exact.
Proof.
Each exact sequence $0\to X\to Y\to Z\to 0$ in $\mathcal{A}$ gives rise to an exact triangle $X\to Y\to Z\to X[1]$ in $\mathbf{D}(\mathcal{A})$. The right derived functor $\mathbf{R}P$ is exact and $\mathbf{D}(\mathcal{U})$ is a triangulated subcategory which is closed under direct summands. Then it follows that $\mathcal{A}_{\mathcal{U}}$ is thick. We have already seen that thickness of $\mathcal{A}_{\mathcal{U}}$ for each $\mathcal{U}$ implies the exactness of the corresponding support function. ∎
Cohomologically stable subcategories
The following definition is the derived analogue of Definition 3.1.
Definition 4.3.
We call a thick subcategory $\mathcal{C}\subseteq\mathcal{A}$ cohomologically stable if
  1. (1)
    for every family $(X_{\alpha})$ of objects in $\mathcal{C}$ the injective envelope of $\coprod_{\alpha}X_{\alpha}$ belongs to $\mathcal{C}$, and
  2. (2)
    for every exact sequence $X^{0}\to X^{1}\to X^{2}\to\cdots$ with all $X^{n}$ in $\mathcal{C}$ the kernel of $X^{0}\to X^{1}$ belongs to $\mathcal{C}$.
The cohomologically stable subcategories are partially ordered by inclusion and closed under arbitrary intersections; so they form a complete lattice. When $\mathcal{A}$ is spectral, then its cohomologically stable subcategories are precisely its localising subcategories. For an object $X\in\mathcal{A}$ let $\langle X\rangle_{\mathrm{ex}}$ denote the smallest cohomologically stable subcategory of $\mathcal{A}$ that contains $X$.
Lemma 4.4.
For $\mathcal{U}\in\mathbf{S}(\mathcal{A})$ the subcategory $\mathcal{A}_{\mathcal{U}}$ is cohomologically stable.
Proof.
The category $\mathcal{A}_{\mathcal{U}}$ is thick by Lemma 4.2. Let $(X_{\alpha})$ be a family of objects in $\mathcal{A}_{\mathcal{U}}$ and choose injective envelopes $X_{\alpha}\to I_{\alpha}$. Then $I_{\alpha}\in\operatorname{Inj}_{\mathcal{U}}\mathcal{A}$ for all $\alpha$, and therefore $E(\coprod_{\alpha}X_{\alpha})\cong E(\coprod_{\alpha}I_{\alpha})$ belongs to $\operatorname{Inj}_{\mathcal{U}}\mathcal{A}$. Now let
$X^{0}\to X^{1}\to X^{2}\to\cdots$
be exact with all $X^{n}$ in $\mathcal{A}_{\mathcal{U}}$. We set $Z^{n}=\operatorname{Ker}(X^{n}\to X^{n+1})$ and need to show that $Z^{0}$ belongs to $\mathcal{A}_{\mathcal{U}}$. Choose an injective envelope $X^{0}\to I^{0}$. The cokernel $C=\operatorname{Coker}(Z^{0}\to I^{0})$ fits into an exact sequence $0\to Z^{1}\to C\to C^{\prime}\to 0$ with $C^{\prime}=\operatorname{Coker}(X^{0}\to I^{0})$ in $\mathcal{A}_{\mathcal{U}}$. We form the pushout
[…]
and the object $\bar{X}^{1}$ belongs to $\mathcal{A}_{\mathcal{U}}$. Thus we obtain an exact sequence
$I^{0}\to\bar{X}^{1}\to X^{2}\to X^{3}\to\cdots$
with all terms in $\mathcal{A}_{\mathcal{U}}$ and $Z^{0}=\operatorname{Ker}(I^{0}\to\bar{X}^{1})$. Proceeding by induction we obtain an exact sequence
$I^{0}\to I^{1}\to I^{2}\to I^{3}\to\cdots$
with all terms in $\operatorname{Inj}_{\mathcal{U}}\mathcal{A}$ and $Z^{0}=\operatorname{Ker}(I^{0}\to I^{1})$. Thus $Z^{0}$ belongs to $\mathcal{A}_{\mathcal{U}}$. ∎
Theorem 4.5.
Let $\mathcal{A}$ be a Grothendieck category. The assignment $\mathcal{C}\mapsto\operatorname{supp}_{\mathrm{ex}}(\mathcal{C})$ induces a lattice isomorphism
$\{\mathcal{C}\subseteq\mathcal{A}\mid\mathcal{C}\text{ cohomologically stable}\}\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{S}(\mathcal{A}).$
The inverse map is given by $\mathcal{U}\mapsto\mathcal{A}_{\mathcal{U}}$.
Proof.
Lemma 4.4 shows that the assignment $\mathcal{U}\mapsto\mathcal{A}_{\mathcal{U}}$ is well defined. Using the description of support in Lemma 4.1, it is clear that $\operatorname{supp}_{\mathrm{ex}}(\mathcal{A}_{\mathcal{U}})=\mathcal{U}$. Now let $\mathcal{C}\subseteq\mathcal{A}$ be a cohomologically stable subcategory and set $\mathcal{U}=\operatorname{supp}_{\mathrm{ex}}(\mathcal{C})$. We claim that $\operatorname{Inj}_{\mathcal{U}}\mathcal{A}\subseteq\mathcal{C}$. When $X\to I$ is a minimal injective resolution of an object in $X\in\mathcal{C}$, then the definition of cohomologically stable implies that $I^{n}\in\mathcal{C}$ for all $n\in\mathbb{Z}$. Moreover, cohomologically stable implies that any object $I\in\operatorname{Inj}\mathcal{A}$ belongs to $\mathcal{C}$ when $P(I)$ arises as a direct summand of a coproduct of objects of the form $P(I^{n})$. Here we use again that $P$ preserves coproducts. This yields the claim, and the equality $\mathcal{A}_{\mathcal{U}}=\mathcal{C}$ then follows. ∎
Remark 4.6.
When $\mathcal{A}$ is locally noetherian, then a cohomologically stable subcategory is closed under all coproducts. This follows from the fact that any coproduct of injectives is again injective. Moreover, in this case $\mathbf{S}(\mathcal{A})$ identifies with the lattice of subsets of $\operatorname{Sp}\mathcal{A}$; see Example 2.6.
Remark 4.7.
In this generality Theorem 4.5 seems to be new. For some special cases it was already known that subcategories of $\mathcal{A}$ are determined by subsets of $\operatorname{Sp}\mathcal{A}$.
(1) When $\mathcal{A}$ equals the category of modules over a commutative noetherian ring $\Lambda$, then $\mathbf{S}(\mathcal{A})$ identifies with the lattice of subsets of the prime ideal spectrum of $\Lambda$; see Example 2.6. In this case the localising subcategories of $\mathcal{A}$ are cohomologically stable and correspond to the specialisation closed subsets of $\operatorname{Spec}\Lambda$. This goes back to work of Gabriel in [10]; it was extended by Takahashi to a correspondence for all cohomologically stable subcategories in [35] and more recently in [25], using Neeman’s classification of localising subcategories of the derived category $\mathbf{D}(\mathcal{A})$ from [27]. A generalisation of Gabriel’s correspondence for locally noetherian Grothendieck categories was established in [36].
(2) For a locally coherent Grothendieck category $\mathcal{A}$, the localising subcategories of finite type are classified in terms of $\operatorname{Sp}\mathcal{A}$; see [16, 21]. This is somewhat orthogonal to Gabriel’s correspondence. More precisely, a localising subcategory $\mathcal{C}\subseteq\mathcal{A}$ is of finite type when the corresponding subcategory of torsion-free objects
$\mathcal{C}^{\perp}=\{X\in\mathcal{A}\mid\operatorname{Hom}(C,X)=0\text{ for all }C\in\mathcal{C}\}$
is closed under filtered colimits. The subcategory $\mathcal{C}^{\perp}$ is an essentially closed subcategory, and both $\mathcal{C}^{\perp}\cap\operatorname{Sp}\mathcal{A}$ and $P(\mathcal{C}^{\perp})\in\mathbf{S}(\mathcal{A})$ determine $\mathcal{C}$, while $\mathcal{C}$ need not be cohomologically stable; cf. Theorem 3.18.
Support data
[…]
Definition 4.8.
An exact support datum on a Grothendieck category $\mathcal{A}$ is a pair $(T,\sigma)$ consisting of a topological space $T$ and a map $\sigma\colon\operatorname{Ob}\mathcal{A}\to\operatorname{Clop}(T)$ such that
  1. (E00)
    $\sigma(X)=\varnothing$ for $X=0$,
  2. (E1111)
    $\sigma(X)=T$ for every generator $X$ of $\mathcal{A}$,
[…]
    $\sigma(X)\cup\sigma(Y)=\sigma(X\oplus Y)$ for all objects $X,Y$ in $\mathcal{A}$, and
[…]
    $\sigma(X)\cap\sigma(Y)=\sigma(Z)$ for all objects $X,Y,Z$ in $\mathcal{A}$ with $\langle X\rangle_{\mathrm{ex}}\cap\langle Y\rangle_{\mathrm{ex}}=\langle Z\rangle_{\mathrm{ex}}$.
A morphism $(T,\sigma)\to(T^{\prime},\sigma^{\prime})$ is given by a continuous map $f\colon T\to T^{\prime}$ such that $\sigma(X)=f^{-1}(\sigma^{\prime}(X))$ for all $X\in\mathcal{A}$.
The following is the exact analogue of Theorem 3.10; it has the same proof.
Theorem 4.9.
The pair $(\operatorname{Spc}\mathcal{A},\operatorname{supp}_{\mathrm{ex}})$ is an exact support datum on $\mathcal{A}$. Moreover, for any exact support datum $(T,\sigma)$ there is a unique continuous map $f\colon T\to\operatorname{Spc}\mathcal{A}$ such that $\sigma(X)=f^{-1}(\operatorname{supp}_{\mathrm{ex}}(X))$ for all $X\in\mathcal{A}$. Explicitly, the map $f$ is given by the adjunction (2.3) and therefore
$f(p)=\{\langle X\rangle\in\mathbf{S}(\mathcal{A})\mid p\not\in\sigma(X)\}\qquad\text{for }p\in T.\qed$
The computation of $\operatorname{supp}_{\mathrm{ex}}$ in Lemma 4.1 shows that for each object $X\in\mathcal{A}$ we have an inclusion
$\operatorname{supp}(X)\subseteq\operatorname{supp}_{\mathrm{ex}}(X).$
There is an equality when $\mathcal{A}$ is spectral, but small examples show that it is a proper inclusion in general.
Example 4.10.
Let $\mathcal{A}=\operatorname{Mod}\Lambda$ be the module category of $\Lambda=\left[\begin{smallmatrix}k&k\\ 0&k\end{smallmatrix}\right]$ where $k$ is a field. Then there are two indecomposable injective $\Lambda$-modules and $\operatorname{Sp}\mathcal{A}=\{I_{1},I_{2}\}=\operatorname{Spc}\mathcal{A}$. There is a short exact sequence $0\to S_{1}\to I_{1}\to I_{2}\to 0$ where $S_{1}=\operatorname{soc}(I_{1})$, and therefore
$\operatorname{supp}(S_{1})=\{I_{1}\}\qquad\text{and}\qquad\operatorname{supp}_{\mathrm{ex}}(S_{1})=\{I_{1},I_{2}\}.$
Stable localising subcategories
Let $\mathcal{A}$ be a Grothendieck category. A localising subcategory is called stable if it is closed under essential extensions, equivalently if it is closed under injective envelopes. The stable localising subcategories are partially ordered by inclusion and closed under all intersections; so they form a complete lattice.
Given an element $\mathcal{U}\in\mathbf{S}(\mathcal{A})$, we have always the inclusion $\mathcal{A}_{\mathcal{U}}\subseteq P^{-1}(\mathcal{U})$. So one may ask when the equality $\mathcal{A}_{\mathcal{U}}=P^{-1}(\mathcal{U})$ holds.
Lemma 4.11.
For a full subcategory $\mathcal{C}\subseteq\mathcal{A}$ the following are equivalent.
  1. (1)
    $\mathcal{C}$ is stable localising.
  2. (2)
    $\mathcal{C}$ is essentially closed and cohomologically stable.
  3. (3)
    There exists $\mathcal{U}\in\mathbf{S}(\mathcal{A})$ such that $\mathcal{C}=\mathcal{A}_{\mathcal{U}}=P^{-1}(\mathcal{U})$.
Proof.
If $\mathcal{C}\subseteq\mathcal{A}$ is stable localising, then $\mathcal{C}=P^{-1}(\mathcal{U})$ for some $\mathcal{U}\in\mathbf{S}(\mathcal{A})$. Also, we have $\mathcal{C}=\mathcal{A}_{\mathcal{U}}$ since $\mathcal{C}$ is cohomologically stable. On the other hand, when $\mathcal{A}_{\mathcal{U}}=P^{-1}(\mathcal{U})$, then this subcategory is clearly stable localising. ∎
Proposition 4.12.
Let $\mathcal{A}$ be a Grothendieck category. The assignment
$\mathcal{C}\longmapsto\operatorname{supp}(\mathcal{C})=P(\mathcal{C})$
induces an embedding
$\{\mathcal{C}\subseteq\mathcal{A}\mid\mathcal{C}\text{ stable localising}\}\hookrightarrow\mathbf{S}(\mathcal{A})$
which preserves arbitrary joins and meets.
Proof.
We apply Lemma 4.11. Each stable localising subcategory is essentially closed. Thus the embedding follows from Theorem 3.3. The stable localising subcategories of $\mathcal{A}$ are closed under arbitrary intersections. So arbitrary meets are preserved. For a family $(\mathcal{C}_{\alpha})_{\alpha\in A}$ of stable localising subcategories with $\mathcal{U}_{\alpha}=P(\mathcal{C}_{\alpha})$ we consider the join $\mathcal{U}=\bigvee_{\alpha}\mathcal{U}_{\alpha}$ and need to show that $P^{-1}(\mathcal{U})$ is a stable localising subcategory. The subcategory is essentially closed. So it suffices to show that for any epimorphism $\phi\colon I\to X$ with $I$ injective and $P(I)\in\mathcal{U}$ we have that $P(X)\in\mathcal{U}$. We begin with the finite case and may assume $A=\{\alpha,\beta\}$. Let $I=I_{\alpha}\oplus I_{\beta}$ with $I_{\alpha}\in\mathcal{C}_{\alpha}$ and $I_{\beta}\in\mathcal{C}_{\beta}$. This decomposition induces an exact sequence $0\to X_{\alpha}\to X\to X_{\beta}\to 0$ where $X_{\alpha}=\phi(I_{\alpha})$ and $X_{\beta}$ is a quotient of $I_{\beta}$. Since $\mathcal{C}_{\alpha}$ and $\mathcal{C}_{\beta}$ are closed under quotients, it follows that $P(X)\in\mathcal{U}$. Writing $A$ as directed union of its finite subsets, we may assume that $\mathcal{U}=\bigvee_{\alpha}\mathcal{U}_{\alpha}$ is directed. Then $I=\sum_{\alpha}I_{\alpha}$ is the directed union of injectives $I_{\alpha}\in\mathcal{C}_{\alpha}$ by Lemma 2.2. This implies $X=\sum_{\alpha}X_{\alpha}$, where $X_{\alpha}=\phi(I_{\alpha})$. We have $P(X)=\sum_{\alpha}P(X_{\alpha})$, and this belongs to $\mathcal{U}$ since $P(X_{\alpha})\in\mathcal{U}_{\alpha}$ for all $\alpha$. ∎
Corollary 4.13.
The stable localising subcategories of a Grothendieck category form a frame and a coframe.
Proof.
This follows from the above theorem since $\mathbf{S}(\mathcal{A})$ is a frame and a coframe, by Propositions 2.3 and 2.4. ∎
Example 4.14.
Let $\mathcal{A}=\operatorname{Mod}\Lambda$ be the module category of a ring $\Lambda$. When $\Lambda$ is commutative noetherian, then all localising subcategories are stable. This fails when $\Lambda$ is not commutative. For $\Lambda=\left[\begin{smallmatrix}k&k\\ 0&k\end{smallmatrix}\right]$ and $k$ a field, there are two proper localising subcategories and one of them is not stable.
Derived categories
For a Grothendieck category $\mathcal{A}$ we have alreday considered the right derived functor of the canonical functor $P\colon\mathcal{A}\to\operatorname{Spec}\mathcal{A}$. The restriction $\operatorname{Inj}\mathcal{A}\to(\operatorname{Inj}\mathcal{A})/\operatorname{Rad}(\operatorname{Inj}\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Spec}\mathcal{A}$ induces an exact functor
$\mathbf{K}(\operatorname{Inj}\mathcal{A})\longrightarrow\mathbf{K}(\operatorname{Spec}\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\mathbf{D}(\operatorname{Spec}\mathcal{A})$
where $\mathbf{K}(\operatorname{Inj}\mathcal{A})$ denotes the category of complexes in $\operatorname{Inj}\mathcal{A}$ up to homotopy. Its computation for any $X\in\mathbf{K}(\operatorname{Inj}\mathcal{A})$ is very explicit. By removing non-zero contractible summands we may assume that the complex $X$ is homotopically minimal so that all differentials are radical morphisms; see [24, Proposition 4.3.18]. Then $P(X)$ has zero differentials.
For example, when $\mathcal{A}$ is locally noetherian the triangulated category $\mathbf{K}(\operatorname{Inj}\mathcal{A})$ is compactly generated with category of compacts equivalent to $\mathbf{D}^{\mathrm{b}}(\operatorname{Noeth}\mathcal{A})$, where $\operatorname{Noeth}\mathcal{A}$ denotes the full subcategory of noetherian objects in $\mathcal{A}$; see [23]. Moreover, in this case the functor $\mathbf{K}(\operatorname{Inj}\mathcal{A})\to\mathbf{D}(\operatorname{Spec}\mathcal{A})$ preserves all coproducts.
Given a triangulated category $\mathcal{T}$ with coproducts, let $\mathbf{L}(\mathcal{T})$ denote the lattice of localising subcategories. The inclusion $i\colon\operatorname{Spec}\mathcal{A}\to\mathbf{D}(\operatorname{Spec}\mathcal{A})$ induces a lattice isomorphism
$\mathbf{L}(\mathbf{D}(\operatorname{Spec}\mathcal{A}))\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\mathbf{L}(\operatorname{Spec}\mathcal{A})$
by taking $\mathcal{U}\subseteq\mathbf{D}(\operatorname{Spec}\mathcal{A})$ to $i^{-1}(\mathcal{U})$. The inverse map sends $\mathcal{V}\subseteq\operatorname{Spec}\mathcal{A}$ to $\mathbf{D}(\mathcal{V})$ where we identify
$\mathbf{D}(\mathcal{V})=\{X\in\mathbf{D}(\operatorname{Spec}\mathcal{A})\mid H^{n}(X)\in\mathcal{V}\text{ for all }n\in\mathbb{Z}\}.$
This yields a notion of support for objects $X\in\mathbf{K}(\operatorname{Inj}\mathcal{A})$ with $\operatorname{supp}(X)\subseteq\operatorname{Spc}\mathcal{A}$, since $\mathbf{L}(\operatorname{Spec}\mathcal{A})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Clop}(\operatorname{Spc}\mathcal{A})$. Details and properties of this are left to the interested reader.
5. Exactly definable categories
There are interesting classes of rings where the Boolean spectrum of its module category is a finite set with the discrete topology. For instance this includes all artinian rings. In such a case it is more interesting to study the pure-injective modules and the corresponding pure spectral category, as the class of injective modules is rather small. The notion of purity for modules goes back to Cohn [3] and was later generalised by Crawley-Boevey in the context of locally finitely presented additive categories [4]. A key idea is to identify the relevant additive category $\mathcal{C}$ with the full subcategory of fp-injective objects in a much larger locally coherent Grothendieck category $\mathbf{P}(\mathcal{C})$, the purity category of $\mathcal{C}$. Following [22] such categories $\mathcal{C}$ are called exactly definable, because they identify with the category of exact functors $\mathcal{A}\to\mathrm{Ab}$ from a small abelian category $\mathcal{A}$ to the category of abelian groups.
Exactly definable categories
Let $\mathcal{A}$ be a Grothendieck category that is locally coherent. Thus $\mathcal{A}$ has a generating set of finitely presented objects and the full subcategory $\operatorname{fp}\mathcal{A}$ of finitely presented objects is abelian. An object $X\in\mathcal{A}$ is fp-injective if $\operatorname{Ext}^{1}(F,X)=0$ for all $F\in\operatorname{fp}\mathcal{A}$. Let $\mathcal{C}:=\operatorname{Fpinj}\mathcal{A}$ denote the full subcategory of fp-injective objects in $\mathcal{A}$. This is an exact subcategory of $\mathcal{A}$ that is closed under filtered colimits and products. Note that the exact structure of $\mathcal{C}$ is intrinsic: a sequence $0\to X\to Y\to Z\to 0$ in $\mathcal{C}$ is exact if and only if it is pure-exact, so a filtered colimit of split exact sequences. Moreover, $\mathcal{C}$ admits pure-injective envelopes, which identify with the injective envelopes in the ambient category $\mathcal{A}$.
Definition 5.1.
A category of the form $\operatorname{Fpinj}\mathcal{A}$ for a locally coherent category $\mathcal{A}$ is called exactly definable; it has filtered colimits, products and an intrinsic exact structure given by the pure-exact sequences.
Let $\mathcal{C}$ be an exactly definable category, given by a locally coherent category $\mathcal{A}$. Then $\mathcal{A}$ is uniqueley determined by $\mathcal{C}$, because the category of pure-injective objects in $\mathcal{C}$ identifies with the category of injectve objects in $\mathcal{A}$, and any abelian category with enough injective objects is uniquely determined by its category of injective objects. We write $\mathbf{P}(\mathcal{C})$ for $\mathcal{A}$ and call it the purity category of $\mathcal{C}$.
Example 5.2.
Let $\Lambda$ be a ring and let $\operatorname{mod}\Lambda$ denote its category of finitely presented modules. The category $(\operatorname{mod}(\Lambda^{\mathrm{op}}),\mathrm{Ab})$ of additive functors $\operatorname{mod}(\Lambda^{\mathrm{op}})\to\mathrm{Ab}$ is a locally coherent Grothendieck category, and the assignment
$X\longmapsto X\otimes_{\Lambda}-|_{\operatorname{mod}(\Lambda^{\mathrm{op}})}$
identifies $\operatorname{Mod}\Lambda$ with the full subcategory of fp-injective objects in $(\operatorname{mod}(\Lambda^{\mathrm{op}}),\mathrm{Ab})$. Thus $\operatorname{Mod}\Lambda$ is an exactly definable category and
$\mathbf{P}(\operatorname{Mod}\Lambda)=(\operatorname{mod}(\Lambda^{\mathrm{op}}),\mathrm{Ab}).$
This goes back to Gruson and Jensen [15]; see also [4, §3] or [24, 12.4].
Definable subcategories
Fix an exactly definable category $\mathcal{C}$, given by a locally coherent category $\mathcal{A}$. Following Crawley-Boevey [5] a full subcategory $\mathcal{D}\subseteq\mathcal{C}$ is called definable if it is closed under filtered colimits, products, and pure subobjects. The terminology is justified by the fact that definable subcategories of $\mathcal{C}$ correspond to Serre subcategories of $\operatorname{fp}\mathcal{A}$: we have
$\mathcal{D}=\{X\in\mathcal{C}\mid\operatorname{Hom}(F,X)=0\text{ for all }F\in\mathcal{S}\}$
for some Serre subcategory $\mathcal{S}\subseteq\operatorname{fp}\mathcal{A}$; see [24, 12.2] for details. From the above description it follows that $\mathcal{D}$ is a thick subcategory. Moreover, $\mathcal{D}$ is closed under pure-injective envelopes; see [24, Proposition 12.2.8]. We generalise the definition of a definable subcategory in two directions, following the Definitions 3.1 and 4.3.
Definition 5.3.
A full subcategory $\mathcal{D}\subseteq\mathcal{C}$ is called pure-essentially closed if it is closed under arbitrary coproducts, pure subobjects, and pure-injective envelopes.
Definition 5.4.
A thick subcategory $\mathcal{D}\subseteq\mathcal{C}$ is called pure-cohomologically stable if
  1. (1)
    for every family $(X_{\alpha})$ of objects in $\mathcal{D}$ the pure-injective envelope of $\coprod_{\alpha}X_{\alpha}$ belongs to $\mathcal{D}$, and
  2. (2)
    for every pure-exact sequence $X^{0}\to X^{1}\to X^{2}\to\cdots$ with all $X^{n}$ in $\mathcal{D}$ the kernel of $X^{0}\to X^{1}$ belongs to $\mathcal{D}$.
For an exact category we recall that a sequence of composable morphisms $(\phi^{n})$ is exact if each $\phi^{n}$ is the composite $\phi^{n}=\iota_{n}\pi_{n}$ of an admissible epimorphism (deflation) followed by an admissible monomorphism (inflation) so that each pair $(\iota_{n},\pi_{n+1})$ is a conflation.
Lemma 5.5.
Any definable subcategory of an exactly definable category is pure-essentially closed and pure-cohomologically stable.
Proof.
This follows from the defining properties of a definable subcategory and the additional properties that have already been mentioned. Specifically, one uses that a definable subcategory is thick and closed under pure-injective envelopes. ∎
The pure-essentially closed and the pure-cohomologically stable subcategories are partially ordered by inclusion and closed under arbitrary intersections; so they form complete lattices.
The Ziegler topology
The class of pure-injective modules has been studied in seminal work of Ziegler on the model theory of modules [37]. In particular, he introduced for any ring a topology on the set of isoclasses of indecomposable pure-injective modules.
Let $\operatorname{Ind}\mathcal{C}$ denote the set of isoclasses of indecomposable pure-injective objects in $\mathcal{C}$; this identifies with $\operatorname{Sp}\mathbf{P}(\mathcal{C})$. The assignment
(5.1) $\mathcal{D}\longmapsto\mathcal{D}\cap\operatorname{Ind}\mathcal{C}$
identifies the definable subcategories of $\mathcal{C}$ with the Ziegler closed subsets of $\operatorname{Ind}\mathcal{C}$. The inverse map sends a Ziegler closed subset $\mathcal{U}$ to the definable subcategory generated by $\mathcal{U}$, which consists of all pure subobjects of products of objects in $\mathcal{U}$. This correspondence is due to Crawley-Boevey for module categories [5, 2.5] and we refer to [24, 12.2] for the general version.
The Boolean spectrum
Let $\operatorname{PInj}\mathcal{C}$ denote the full subcategory of pure-injective objects in $\mathcal{C}$; it identifies with $\operatorname{Inj}\mathbf{P}(\mathcal{C})$ via the embedding $\mathcal{C}\to\mathbf{P}(\mathcal{C})$. The pure spectral category is the quotient
$\operatorname{PSpec}\mathcal{C}:=(\operatorname{PInj}\mathcal{C})/\operatorname{Rad}(\operatorname{PInj}\mathcal{C})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\operatorname{Spec}\mathbf{P}(\mathcal{C})$
which is a spectral Grothendieck category because of the equivalence from (2.1). We consider the lattice of localising subcategories
$\mathbf{L}(\operatorname{PSpec}\mathcal{C})\xrightarrow{\raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}}\mathbf{S}(\mathbf{P}(\mathcal{C})).$
This is a Boolean lattice by Proposition 2.8 and its spectrum
$\operatorname{PSpc}\mathcal{C}:=\operatorname{Spec}(\mathbf{L}(\operatorname{PSpec}\mathcal{C}))$
can be written as the disjoint union of a discrete and a continuous part as in (2.5). This yields a canonical embedding $\operatorname{Ind}\mathcal{C}\hookrightarrow\operatorname{PSpc}\mathcal{C}$ which identifies $\operatorname{Ind}\mathcal{C}$ with the discrete part. We endow $\operatorname{Ind}\mathcal{C}$ with the Ziegler topology and write $\operatorname{Cl}(\operatorname{Ind}\mathcal{C})$ for its lattice of closed subsets.
In the following we provide two possible extensions of Crawley-Boevey’s correspondence (5.1) for definable subcategories, which amounts to a lattice isomorphism
$\{\mathcal{D}\subseteq\mathcal{C}\mid\mathcal{D}\text{ definable}\}\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\operatorname{Cl}(\operatorname{Ind}\mathcal{C}).$
We view $\mathcal{C}$ as a full subcategory of the purity category $\mathbf{P}(\mathcal{C})$ so that $\operatorname{supp}(X)$ and $\operatorname{supp}_{\mathrm{ex}}(X)$ for objects $X\in\mathcal{C}$ are defined via (3.1) and (4.1). Moreover, these are viewed as subsets of $\operatorname{PSpc}\mathcal{C}$ via the isomorphism
[…]
Theorem 5.6.
Let $\mathcal{C}$ be an exactly definable category. The assignment $\mathcal{D}\mapsto\operatorname{supp}(\mathcal{D})$ induces a lattice isomorphism
$\{\mathcal{D}\subseteq\mathcal{C}\mid\mathcal{D}\text{ pure-essentially closed}\}\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\operatorname{Clop}(\operatorname{PSpc}\mathcal{C}).$
Moreover, the embedding $\operatorname{Ind}\mathcal{C}\hookrightarrow\operatorname{PSpc}\mathcal{C}$ identifies $\mathcal{D}\cap\operatorname{Ind}\mathcal{C}$ with $\operatorname{supp}(\mathcal{D})\cap\operatorname{Ind}\mathcal{C}$ when $\mathcal{D}\subseteq\mathcal{C}$ is definable.
Proof.
We view $\mathcal{C}\subseteq\mathbf{P}(\mathcal{C})$ as a full subcategory. Taking an essentially closed subcategory $\mathcal{U}\subseteq\mathbf{P}(\mathcal{C})$ to $\mathcal{U}\cap\mathcal{C}$ induces a bijection between the essentially closed subcategories of $\mathbf{P}(\mathcal{C})$ and the pure-essentially closed subcategories of $\mathcal{C}$. This follows from the fact that any essentially closed subcategory $\mathcal{U}\subseteq\mathbf{P}(\mathcal{C})$ is determined by $\mathcal{U}\cap\operatorname{Inj}\mathbf{P}(\mathcal{C})$, keeping in mind that $\operatorname{PInj}\mathcal{C}=\operatorname{Inj}\mathbf{P}(\mathcal{C})$. Given this correspondence the isomorphism is obtained by composing the isomorphisms from Theorem 3.3 and (5.2). The additional assertion about definable subcategories follows by unraveling the construction of the embedding $\operatorname{Ind}\mathcal{C}\hookrightarrow\operatorname{PSpc}\mathcal{C}$. ∎
Theorem 5.7.
Let $\mathcal{C}$ be an exactly definable category. The assignment $\mathcal{D}\mapsto\operatorname{supp}_{\mathrm{ex}}(\mathcal{D})$ induces a lattice isomorphism
$\{\mathcal{D}\subseteq\mathcal{C}\mid\mathcal{D}\text{ pure-cohomologically stable}\}\xrightarrow{\ \raisebox{-1.72218pt}[0.0pt][0.0pt]{$\scriptstyle{\sim}$}\ }\operatorname{Clop}(\operatorname{PSpc}\mathcal{C}).$
Moreover, the embedding $\operatorname{Ind}\mathcal{C}\hookrightarrow\operatorname{PSpc}\mathcal{C}$ identifies $\mathcal{D}\cap\operatorname{Ind}\mathcal{C}$ with $\operatorname{supp}_{\mathrm{ex}}(\mathcal{D})\cap\operatorname{Ind}\mathcal{C}$ when $\mathcal{D}\subseteq\mathcal{C}$ is definable.
Proof.
Adapt the proof of Theorem 5.6 by using Theorem 4.5 instead of Theorem 3.3. ∎
Remark 5.8.
For every definable subcategory $\mathcal{D}\subseteq\mathcal{C}$ we have $\operatorname{supp}(\mathcal{D})=\operatorname{supp}_{\mathrm{ex}}(\mathcal{D})$.
Module categories
Let $\Lambda$ be a ring and consider its module category $\mathcal{C}=\operatorname{Mod}\Lambda$ which is exactly definable. We use the obvious notation and set $\operatorname{PInj}\Lambda=\operatorname{PInj}\mathcal{C}$, $\operatorname{Ind}\Lambda=\operatorname{Ind}\mathcal{C}$, and $\operatorname{PSpc}\Lambda=\operatorname{PSpc}\mathcal{C}$. The books of Prest [28] and also Jensen and Lenzing [17] discuss the basic structure of the pure spectral category
$\operatorname{PSpec}\Lambda=(\operatorname{PInj}\Lambda)/\operatorname{Rad}(\operatorname{PInj}\Lambda).$
For instance, it is shown that any pure-injective module $X$ admits an essentially unique decomposition $X=X_{\mathrm{d}}\oplus X_{\mathrm{c}}$ into a discrete and a continuous part [17, Corollary 8.28]. This reflects the decomposition
$\operatorname{PSpec}\Lambda=(\operatorname{PSpec}\Lambda)_{\mathrm{d}}\times(\operatorname{PSpec}\Lambda)_{\mathrm{c}}$
of the spectral category. For a more detailed treatment of decompositions of pure-injective modules we refer to [9]. Unfortunately, not much seems to be known about the continuous part of the spectral category, except that one knows for various classes of rings when the continuous part vanishes. On the other hand, superdecomposable modules have been constructed more or less explicitly for several rings, but these constructions do not provide much insight into the ways they decompose; see [31] for a survey. In any case, it is clear from Proposition 3.6 that the Boolean lattice $\mathbf{L}(\operatorname{PSpec}\Lambda)$ controls the decompositions of pure-injective $\Lambda$-modules.
Proposition 5.9.
We have $X=X_{\mathrm{d}}$ for every pure-injective $\Lambda$-module $X$ if and only if the canonical embedding $\operatorname{Ind}\Lambda\hookrightarrow\operatorname{PSpc}\Lambda$ is a bijection.
Proof.
This follows from the above decomposition of the pure spectral category $\operatorname{PSpec}\Lambda$ and the corresponding decomposition of the Boolean spectrum (2.5). ∎
The following result complements our extension of Crawley-Boevey’s correspondence for definable subcategories.
Proposition 5.10.
For an Artin algebra $\Lambda$ the following are equivalent.
  1. (1)
    The embedding $\operatorname{Cl}(\operatorname{Ind}\Lambda)\to\operatorname{Clop}(\operatorname{PSpc}\Lambda)$ is a bijection.
  2. (2)
    Every pure-essentially closed subcategory of $\operatorname{Mod}\Lambda$ is definable.
  3. (3)
    Every pure-cohomologically stable subcategory of $\operatorname{Mod}\Lambda$ is definable.
  4. (4)
    The Ziegler spectrum $\operatorname{Ind}\Lambda$ is a discrete space.
  5. (5)
    The algebra $\Lambda$ is of finite representation type.
Proof.
(1) $\Leftrightarrow$ (2): This follows from Theorem 5.6.
(1) $\Leftrightarrow$ (3): This follows from Theorem 5.7.
(2) $\Rightarrow$ (4): Every subset of $\operatorname{PSpc}\Lambda$ contained in $\operatorname{Ind}\Lambda$ is clopen, by Lemma 2.5. It follows from Theorem 5.6 that every subset of $\operatorname{Ind}\Lambda$ is Ziegler closed.
(4) $\Rightarrow$ (5): The space $\operatorname{Ind}\Lambda$ is finite when it is discrete, because $\operatorname{Ind}\Lambda$ is quasi-compact. For an Artin algebra, every indecomposable finitely presented module is pure-injective, so $\Lambda$ is of finite representation type.
(5) $\Rightarrow$ (2): When $\Lambda$ is of finite representation type, then every $\Lambda$-module is endofinite, and in particular pure-injective [24, Theorem 13.2.10]. Let $\mathcal{D}$ be a pure-essentially closed subcategory. Then it follows that $\mathcal{D}$ is closed under pure subobjects and pure quotients. In particular, $\mathcal{D}$ is closed under filtered colimits because any filtered colimit of objects in $\mathcal{D}$ is a pure quotient of a coproduct of objects in $\mathcal{D}$. Finite representation type implies that every $\Lambda$-module decomposes into a coproduct of indecomposables, and every product of objects in $\mathcal{D}$ decomposes into indecomposables from $\mathcal{D}$, because only finitely many indecomposables are involved. This follows again from endofiniteness. Thus $\mathcal{D}$ is closed under products. ∎
We close our discussion of module categories with some remarks.
Remark 5.11.
Beyond finite representation type the correspondences for pure-essentially closed and pure-cohomologically stable subcategories (Theorems 5.6 and 5.7) seem to be new. In particular, Proposition 5.10 shows that the extension of the Ziegler spectrum yields additional information whenever the ring is not of finite representation type. This is illustrated by looking at the category of abelian groups $\mathrm{Ab}=\operatorname{Mod}\mathbb{Z}$. Consider the Prüfer group $\mathbb{Z}(p^{\infty})$ given by a prime $p$. The direct sums of copies of $\mathbb{Z}(p^{\infty})$ form a subcategory of $\mathrm{Ab}$ which is pure-essentially closed and pure-cohomologically stable, but not definable, because it is not closed under products. More precisely, the rationals $\mathbb{Q}$ arise as direct summand of a product of copies of $\mathbb{Z}(p^{\infty})$, and then $\{\mathbb{Z}(p^{\infty}),\mathbb{Q}\}$ is the Ziegler closure of $\mathbb{Z}(p^{\infty})$.
The Ziegler spectrum plus all definable subcategories are well understood for various classes of rings, including the rings of finite representation type, commutative Dedekind domains, or tame hereditary algebras; see [29] for a survey. For these rings it seems reasonable to expect similar descriptions for the pure-essentially closed and the pure-cohomologically stable subcategories. Note that in these examples $\operatorname{Ind}\Lambda=\operatorname{PSpc}\Lambda$; so the relevant subcategories are parameterised by all subsets of $\operatorname{Ind}\Lambda$.
Acknowledgements
It is a pleasure to thank Greg Stevenson for several discussions and helpful comments on this work. The work was supported by the Deutsche Forschungsgemeinschaft (Project-ID 491392403 – TRR 358).